OpenStudy (anonymous):

laplace transform of cos^3a*t

OpenStudy (anonymous):

best proof i could find for manipulating cos^3(at), which is the only tricky part here: We can write, cos(3x)=cos(2x+x) =cos(2x)cos(x) - sin(2x)sin(x) ={2cos^2(x) - 1}cos(x) - 2sin(x)cos(x)sin(x) =2cos^3(x) - cos(x) - 2cos(x).{1-cos^2(x)} =2cos^3(x) - cos(x) - 2cos(x) + 2cos^3(x) =4cos^3(x) - 3cos(x) Now, 4cos^3(x) = cos(3x)+3cos(x) so, cos^3(x)= 1/4[cos(3x)+3cos(x)] from yahoo answers: http://au.answers.yahoo.com/question/index?qid=20110404031547AABJeNW laplace transform of that should be trivial :)

OpenStudy (anonymous):

or you could just derive it \[F(s)= \int\limits_{0}^{\infty}e^{-st}f(x)dx\]

OpenStudy (anonymous):

implying you have all day :)

OpenStudy (callisto):

*Little correction \[F(s)= \int\limits_{0}^{\infty}e^{-st}f(t)dt\]